OpenStudy (anonymous):

Simplify −3(d − 3) + 5d

OpenStudy (jackellyn):

Distribute first, what do you get? Then add that to 5d (combining like terms only)

Parth (parthkohli):

You can start this by distributing. \( \color{Black}{\Rightarrow (-3 \times d) - (-3 \times 3) + 5d }\)

Parth (parthkohli):

And then combine like terms :)

OpenStudy (anonymous):

I'm Confused

OpenStudy (anonymous):

-3d+5d+9=

OpenStudy (jackellyn):

First you distribute -3(d-3) which gives you -3d+9 and you add that to 5d so -3d+5d+9=?

OpenStudy (anonymous):

−3d + 9 2d + 9 2d − 9 2d − 3

OpenStudy (anonymous):

-3(d-3)+5d= -3d+9+5d= 2d+9

OpenStudy (jackellyn):

Why did you do steps 3 & 4 in your solution?

OpenStudy (anonymous):

Me?

OpenStudy (jackellyn):

Yes

OpenStudy (anonymous):

Those Are The Answer Choices

OpenStudy (jackellyn):

Oh okay, sorry. :)